Tuesday, November 01, 2005

Week of 10-24-05

Specific project tasks completed this week include:

- Measurement data on laser output power stability taken last week was analyzed. It was found that any output power instability that could be the cause of the propagation time shift noise jitter through the crystal that has been observed would be fairly minimal, and below the detection sensitivity of the detectors used last week (ThorLabs). Higher sensitivity detectors were located and new measurements on power stability will be taken (same time scale - every 1 min for 1 hr).

- Cleaned up the lab!

- Midterms.........

0 Comments:

Post a Comment

<< Home